Status Report

SOHO Daily Meeting Minutes for Tuesday, 25 June 2002

By SpaceRef Editor
June 25, 2002
Filed under , ,

* RHESSI target for today is AR 0008; target for tomorrow
is TBD.

* Visitors will come to the EOF on Tuesday 25 June in the afternoon,
on Thursday June 27, and Friday 28 June

* SPWG Friday at 2pm in the EAF. Please get calendar changes
to the SOC as soon as possible.

* SDAC computers (including the archive) will be down for
3 hours today for maintenance.

* Today from 5:00 pm until 7:00 pm, there will be an outage
of the fire alarm system in building 14. Bells will ring
during test and you can ignore them.

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, New Tracking Star

Planned Activities: RSL, VIRGO, New Guide Star (3.0)

Upcoming Operations: 7/1 ACU/CDMU Memory Dumps

Ground Anomalies:
175/1426 Wayward command. D27. 4 minute commanding delay.
176/1200 4 minute late AOS. D16. Complex supervisor crash.
TM recoverable from SSR.
176/1207 3 minutes late commanding. D16. CPA failure.

06/24 DOY 175 – 06/25 DOY 176

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. Various sunspot observations on AR0008.
Active region studies for Wednesday and Thursday.

UVCS: Nominal. 24-hour synoptic today, Cassini conjunction and
mini-synoptic tomorrow.

LASCO: Nominal. C2/C3 synoptics.

EIT: Nominal. 195 CME watch and synoptics.

MDI: Nominal. Full disk magnetograms and dopplergrams.

TRACE: Nominal. Supporting RHESSI.

SpaceRef staff editor.